A missile is launched on the earth with the initial velocity of 100ms-1 at an elevation of 30 degree s.calculate


1 greatest height attained


2 horizontal range


3 time of fright


4 velocity of a missile at time 8 second

1
Expert's answer
2022-05-13T17:03:35-0400

1.


"H_{\max}=\frac{v_0^2\sin^2\theta}{2g}""H_{\max}=\frac{100^2\sin^230^\circ}{2*9.8}=128\:\rm m"

2.


"R=\frac{v_0^2\sin2\theta}{g}""R=\frac{100^2\sin60^\circ}{9.8}=884\:\rm m"

3.


"t=\frac{2v_0\sin\theta}{g}""t=\frac{2*100\sin30^\circ}{9.8}=10.2\:\rm s"

4.


"v=\sqrt{(v_0\cos\theta)^2+(v_0\sin\theta-gt)^2}""v=\sqrt{(100\cos30^\circ)^2+(100\sin30^\circ-9.8*8.0)^2}=91\:\rm m/s"